What is the present worth of a P100 annuity starting at the end of the third year and continuing to the end of the fourth year, if the annual interest rate is 8%?

1
Expert's answer
2022-03-16T14:45:58-0400

The present value is:

"PV= \frac{FV}{(1+{r})^n}"

"PV= \frac{100}{(1+{8\%})^3}+ \frac{100}{(1+{8\%})^4}"

"PV=P153"